OpenStudy (anonymous):

integral of xln(x^2+1) dx from (0,1)

OpenStudy (anonymous):

where did the answer go?

OpenStudy (anonymous):

i made a mistake.

OpenStudy (anonymous):

wait i will give u one .

myininaya (myininaya):

let u=x^2+1

OpenStudy (anonymous):

k

myininaya (myininaya):

the antiderivative of ln(x) is x ln(x)-x+C since the derivative of (x ln(x)-x)'=ln(x)+1-1=ln(x)

OpenStudy (anonymous):

in my answer ad 1+1/4

myininaya (myininaya):

so the antiderivative of ln(u) is u ln(u)-u+C since the derivative of (u ln(u)-u)'=ln(u)+u-u=ln(u)

OpenStudy (anonymous):

so I did that all wrong?

OpenStudy (anonymous):

yes its wrong!

OpenStudy (anonymous):

I ended up with -1/4

OpenStudy (anonymous):

or if u think i ur not find the derevative of 1/x. and c if u can.

OpenStudy (anonymous):

it would be -x^-2

myininaya (myininaya):

what? i think you are getting confused \[\int\limits_{}^{}\frac{1}{u} dx=\ln|u|+C\] \[\int\limits_{}^{}\ln(u) du=u \ln(u)-u+C\] \[(\frac{1}{u})'=\frac{-1}{u^2}\]

OpenStudy (anonymous):

okay so is -1/4 the wrong answer?

OpenStudy (anonymous):

yeh ur right im just tired i need to go to sleep.forgot that ,thanks for the alert.

myininaya (myininaya):

it happens to all of us we all get tired

OpenStudy (anonymous):

i was writing physics , but i promise 2moro morning i will be 5n.